OpenStudy (anonymous):

How do you write (x+5)to the 2nd power in expanded form?

OpenStudy (anonymous):

(x+5)(x+5)

OpenStudy (anonymous):

can you solve that

OpenStudy (anonymous):

(X+25)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

(Xto the second + 25) is that write

OpenStudy (anonymous):

\[\left( x+5 \right)\left( x+5 \right)=x \left( x+5 \right)+5\left( x+5 \right)=?\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

(X to the second + 10)